# Addressing Polarization and Unfairness in Performative Prediction

### v1

@@ 3 Unfairness and Polarization in SDPP
In this work, we study SDPP with different demographic groups, where an ML model $\boldsymbol{\theta}$ is trained to make predictions about individuals from multiple groups distinguished by a sensitive attribute $s\in\mathcal{S}$ (e.g., gender, age, race), whose data distribution changes based on the deployed ML model and such model-dependent distribution shift can be captured by transition mapping T. Suppose individuals from group $s$ follow the identical data distribution $\mathcal{D}_{s}^{(t)}$ at the round $t$ of an iterative algorithm, and let $p_{s}^{(t)}$ be the size of group $s$ as the fraction of entire population at $t$. Then the data distribution of the entire population is $\mathcal{D}^{(t)}=\sum_{s\in\mathcal{S}}p_{s}^{(t)}\mathcal{D}_{s}^{(t)}$ with $\sum_{s\in\mathcal{S}}p_{s}^{(t)}=1$.
@@
    Strategic classification with multiple groups [16, 17]: individuals in applications such as lending, hiring, and admission may manipulate their data based on ML model strategically to increase their chances of receiving favorable decisions, leading to changes in group distribution $\mathcal{D}_{s}^{(t)}$.
@@
    Decision-making systems under user retention dynamics [2, 12, 18]: ML models used in recognition or recommendation systems may attract more users if they experience high accuracy but drive away those with less satisfaction, which may cause group proportion $p_{s}^{(t)}$ to change.
In this paper, we first explore the fairness properties of PS solutions in SDPP, i.e., examining whether $\boldsymbol{\theta}^{\text{PS}}$ in Eqn. (1) have disparate impacts on different demographic groups. Specifically, we consider two fairness metrics: group-wise loss disparity $\triangle_{\mathcal{L}}^{(t)}$ [6, 19] and participation disparity $\triangle_{p}^{(t)}$ [8] measuring the difference of loss $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t)})$ and fraction $p_{s}^{(t)}$ across different groups at round $t$ of an iterative algorithm. In examples with two groups $\mathcal{S}=\{a,b\}$, the unfairness can be quantified as
$\triangle_{\mathcal{L}}^{(t)}:=\left|\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{a}^{(t)})-\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{b}^{(t)})\right|,~{}~{}~{}~{}~{}\triangle_{p}^{(t)}:=\left|p_{a}^{(t)}-p_{b}^{(t)}\right|.$ (4)
@@
Figure 1: Dynamics of the group fraction in Example 3.1 (left) and group-wise accuracy in Example 3.2 (right) under RRM: the system converges to PS solution $(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}})$ that is unfair.
@@
We first show (through two toy examples) that PS solutions $\boldsymbol{\theta}^{\text{PS}}$ in SDPP may have disparate impacts on different groups. Specifically, when finding $\boldsymbol{\theta}^{\text{PS}}$ using iterative algorithms introduced in Section 2, the process may incur severe polarization effects and exhibit unfairness, i.e., certain groups may get marginalized and diminish from the system, and group-wise loss disparity gets exacerbated during the iterative process.
Example 3.1 (Polarization effects of $\boldsymbol{\theta}^{\text{PS}}$).
Consider individuals from two groups $a,b$ whose participation in an ML system depends on their perceived loss, i.e., group fraction $p_{s}^{(t)}$ changes based on the deployed ML model. Suppose individuals from group $a$ and $b$ have fixed data $Z=z_{a}=1$ and $Z=z_{b}=0$, respectively. Their initial group fractions are $p_{a}^{(0)}=0.7,p_{b}^{(0)}=0.3$. Consider conventional RRM with mean squared error $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t)})=\sum_{s\in\{a,b\}}p_{s}^{(t)}(\theta-z_{s})^{2}$ and transition T that satisfies $p_{a}^{(t+1)}=p_{a}^{(t)}+0.1\cdot\left[\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{b}^{(t)})-\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{a}^{(t)})\right],p_{b}^{(t+1)}=1-p_{a}^{(t+1)}$ (i.e., group fraction decreases if the group has higher loss). It is easy to see that $p_{a}^{(t)}$ (resp. $p_{b}^{(t)}$) monotonically increases (resp. decreases) in $t$, and the PS solution solely contains people from group $a$ (see Fig. 1).
Example 3.2 (Exacerbated group-wise loss disparity).
Consider individuals from two groups $a,b$ who are subject to certain ML decisions; each has an initial feature $X\sim\text{Uniform}(0,1)$ and a binary label $Y=\mathbbm{1}(X\geq 0.5)\in\{0,1\}$. Suppose the population has fixed group fractions $p_{a}=0.7,p_{b}=0.3$, but individuals may strategically manipulate their features to increase their chances of receiving favorable ML decisions, i.e., distribution of each group $\mathcal{D}_{s}^{(t)}$ changes based on ML model. Following the individual response model in [20, 16], we suppose individuals from two groups have certain effort budgets $\eta_{a}=0.2,\eta_{b}=0.01$ that can be used to improve their features $X_{s}^{(t+1)}=X_{s}^{(t)}+\eta_{s}\cdot\boldsymbol{\theta}^{(t)}$. Consider RRM iterative algorithm where the model parameter $\boldsymbol{\theta}^{(t)}$ is updated to maximize prediction accuracy $\Pr(\mathbbm{1}(X\geq\boldsymbol{\theta}^{(t)})=Y^{(t)})$. Given an initial model $\boldsymbol{\theta}^{(0)}=0.5$, it is trivial to see that $\boldsymbol{\theta}^{(t)}$ increases over time to be robust against manipulation until converges to PS solution $\boldsymbol{\theta}^{\text{PS}}=0.625$ (see Fig. 1). However, at the PS solution, the accuracy for group $a$ is 1 but is only $0.88$ for group $b$.
4 Fair-PS Solutions and Fairness Mechanisms
@@
Section 3 shows that without fairness consideration, PS solution of SDPP may incur polarization effects and have disparate impacts on different groups, i.e., when finding $\boldsymbol{\theta}^{\text{PS}}$ using iterative algorithms such as RRM, group-wise loss disparity $\triangle_{\mathcal{L}}^{(t)}$ and participation disparity $\triangle_{p}^{(t)}$ may get exacerbated and converge to an unfair state. In this section, we address unfairness issues in SDPP. One straightforward idea is to directly apply the fairness mechanisms at every round of the iterative algorithms introduced in Section 2. However, we will show that although such methods are effective in conventional supervised learning, they can disrupt the stability and the iterative algorithms may no longer converge.
4.1 Fairness mechanisms and Fair-PS solutions
@@
    Fairness via regularization: It adds a regularization/penalty term to the original learning objective function $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})$, which penalizes the violation of fairness [9, 10]. The fair objective function is
@@
    Fairness via sample re-weighting: It adjusts the weights of samples (possibly adversarially) and increases sample weights for disadvantaged groups [11, 12, 13]. One example is distributionally robust optimization (DRO) [18], which minimizes the worst-case loss and the fair objective is
    $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D},\rho):=\max_{\tilde{\mathcal{D}}\in\mathcal{B}(\mathcal{D},r(\rho))}\mathcal{L}(\boldsymbol{\theta};\tilde{\mathcal{D}}),$ (6)
@@
$\mathcal{D}^{\text{PS}}_{\textsf{fair}}=\text{T}(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}};\mathcal{D}^{\text{PS}}_{\textsf{fair}}),~{}~{}~{}~{}\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~{}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{\text{PS}}_{\textsf{fair}},\rho).$
@@ 4.2 Existing designs fail to converge to Fair-PS solutions
Example 4.2 (Group loss variance as penalty term).
Consider two groups $a,b$ with data $Z=(X,Y)$ and fixed group fractions $p_{a}=p_{b}=0.5$. Suppose all samples in group $a$ are $(1,-1)$ and all samples in group $b$ are $(2,1)$. Consider squared loss function $\ell(z;\theta)=(y-h_{\theta}(x))^{2}=(y-\theta x)^{2}$ which is strongly convex and jointly smooth. Under group loss variance penalty $\mathcal{P}(\boldsymbol{\theta};\mathcal{D},\rho)$, we have $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D},\rho)=2.5\theta^{2}-\theta+1+\rho\cdot(2.25\theta^{4}+9\theta^{2}-9\theta^{3})$. When $\rho=0.6$, the second-order derivative $\nabla^{2}_{\theta}\mathcal{L}_{\textsf{fair}}=16.2\theta^{2}-32.4\theta+15.8$ and is negative when $\theta=1$. Because convexity of $\mathcal{L}_{\textsf{fair}}$ is a necessary condition for iterative algorithms such as RRM to converge to a stable solution (by Lemma 2.4), adding group loss variance as a penalty at each round will disrupt the stability and the algorithm fails to find Fair-PS solutions. Appendix C.4 provides an empirical illustration of this.
Example 4.3 (Repeated DRO with $\chi^{2}$-distance metric).
Consider two groups $a,b$ with fixed data $z_{a}=1$ and $z_{b}=-1$ for all samples but group fractions $p_{a}^{(t)}=0.5\cdot(1+\theta^{(t)})$ and $p_{b}^{(t)}=0.5\cdot(1-\theta^{(t)})$. $p_{a}^{(0)}=0.4,p_{b}^{(0)}=0.6$. Consider a mean estimation task where the model parameter ${\theta}^{(t+1)}:=\arg\min_{\boldsymbol{\theta}}\max_{\tilde{\mathcal{D}}\in\mathcal{B}(\mathcal{D}^{(t)},r)}\mathcal{L}({\theta};\tilde{\mathcal{D}})$ is updated using DRO with mean squared error and $\chi^{2}$-distance bound $r=1/6$. Let $q_{a}^{(t)},q_{b}^{(t)}$ denote the group fractions of "worst-case" distribution $\tilde{\mathcal{D}}$. We have $q_{a}^{(0)}=0.6,q_{b}^{(0)}=0.4$ and $\theta^{(1)}=\arg\min_{\theta}q_{a}^{(0)}(1-\theta)^{2}+q_{b}^{(0)}(1+\theta)^{2}=0.2$, which results in $p_{a}^{(1)}=0.6,p_{b}^{(1)}=0.4$. Since $\mathcal{L}_{a}(\theta^{(1)};z_{a})<\mathcal{L}_{b}(\theta^{(1)};z_{b})$, DRO should minimize the risk of the "worst-case" distribution with $q_{a}^{(1)}=0.4,q_{b}^{(1)}=0.6$, i.e., $\theta^{(2)}=\arg\min_{\theta}q_{a}^{(1)}(1-\theta)^{2}+q_{b}^{(1)}(1+\theta)^{2}=-0.2$. Repeating the procedure we will get $p_{a}^{(2)}=0.4,p_{b}^{(2)}=0.6$ and $q_{a}^{(2)}=0.6,q_{b}^{(2)}=0.4$, $\theta^{(3)}=0.2$. It turns out that repeated DRO results in $\theta^{(t)}$ oscillating between $0.2$ and $-0.2$ and it never converges.
It is worth noting that DRO methods have been used in Hashimoto et al. [19] to mitigate group fraction disparity in repeated optimizations. However, it only improves fairness without any convergence guarantees to stable solutions. Peet-Pare et al. [21] repeatedly used DRO to improve fairness under PP settings, it only converges to a PS solution under strict assumptions that are hard to verify. Under milder conditions in Lemma 2.4, it may fail to converge as Example 4.3 illustrated.
@@ 4.3 Novel designs for fairness mechanism
Proposed fair regularization (with and without demographics). Depending on whether sensitive attributes $S$ are accessible during training, we propose two fairness penalty terms, as detailed below.
@@
    Group level fairness penalty: It updates model parameter $\boldsymbol{\theta}^{(t)}$ at round $t$ by the following
    $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho):=\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})+\rho\sum_{s\in\mathcal{S}}p^{(t-1)}_{s}[\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})]^{2}$ (7)
@@
    Sample level fairness penalty without demographics: It updates model parameter $\boldsymbol{\theta}^{(t)}$ at round $t$ as follows, which does not need the knowledge of sensitive attributes.
    $\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho):=\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})+\rho\mathbb{E}_{Z\sim\mathcal{D}^{(t-1)}}\big{[}[\ell(\boldsymbol{\theta};Z)]^{2}\big{]}.$ (8)
Proposed fair sample re-weighting. It updates model parameter $\boldsymbol{\theta}^{(t)}$ at round $t$ as follows.
$\boldsymbol{\theta}^{(t)}=\operatorname{argmin}_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho):=\sum_{s\in\mathcal{S}}q_{s}^{(t-1)}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)}),$ (9)
with $\text{with }~{}\boldsymbol{q}^{(t-1)}=\left[q^{(t-1)}_{s}\right]_{s\in\mathcal{S}}:=\frac{\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}}{\|\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}\|_{1}}\text{ and }~{}\boldsymbol{l}^{(t-1)}:=\left[p_{s}^{(t-1)}\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)})\right]_{s\in\mathcal{S}}$
Unlike DRO methods in [22, 19] that require solving a min-max optimization at the current round, our re-weighting method only adjusts the weights for each group based on the group-wise losses in the previous round, which is more computationally efficient.
Comparison & discussion. Intuitively, compared to original $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})$, all three proposed fair objective functions $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ improves fairness at each round by assigning more weights to disadvantaged groups/samples (i.e., the groups/samples experiencing higher losses) in the upcoming update. Indeed, both sample level fairness penalty and fair sample re-weighting can be regarded as modifications of group level fairness penalty.
Comparing Eqn. (7) and (8), the two penalty terms get more similar when most individual samples in the disadvantaged (resp. advantaged) groups are also similarly disadvantaged (resp. advantaged). This is more likely to happen when each demographic group has a distribution with a small variance. For example, if the distribution of each group is a point mass, Eqn. (7) and (8) are identical.
@@
$\|\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}\|_{1}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)=~{}\sum_{s\in\mathcal{S}}p_{s}^{(t-1)}(1+\rho)\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)})\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)}).$
We can see that the right-hand side will be a multiple of Eqn. (7) if we replace $\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)})$ with $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})$. This means both equations yield the same $\theta^{(t)}$ when the population distribution does not change from $t-2$ to $t-1$, suggesting that the two approaches become more similar when the sensitivity of the transition mapping T is smaller, or equivalently, the distribution shift is milder. Finally, we also discuss designing fair objective functions generally in App. B.4.

### v2

@@ 3 Unfairness and polarization in SDPP
In this work, we study SDPP with different demographic groups, where an ML model $\boldsymbol{\theta}$ is trained to make predictions about individuals from multiple groups distinguished by a sensitive attribute $s\in\mathcal{S}$ (e.g., gender, age, race), whose data distribution changes based on the deployed ML model and such model-induced distribution shift can be captured by transition map T. Suppose individuals from group $s$ follow the identical data distribution $\mathcal{D}_{s}^{(t)}$ at the round $t$ of an iterative algorithm, and let $p_{s}^{(t)}$ be the size of group $s$ as the fraction of entire population at $t$. Then the data distribution of the entire population is $\mathcal{D}^{(t)}=\sum_{s\in\mathcal{S}}p_{s}^{(t)}\mathcal{D}_{s}^{(t)}$ with $\sum_{s\in\mathcal{S}}p_{s}^{(t)}=1$.
@@
    Strategic classification (hardt_strategic_2015; Zhang_2022_ICML): individuals in high-stakes applications such as lending, hiring, and college admission may manipulate their data based on ML model strategically to increase their chances of receiving favorable decisions, leading to changes in group distribution $\mathcal{D}_{s}^{(t)}$.
@@
    Decision-making systems under user retention dynamics (Zhang_2019_Retention; duchi2018learning; pmlr-v80-hashimoto18a): ML models in recognition or recommendation systems may attract more users if they experience high accuracy but drive away those with less satisfaction, causing the group proportion $p_{s}^{(t)}$ to change.
We first explore the fairness properties of PS solutions in SDPP, i.e., examining whether $\boldsymbol{\theta}^{\text{PS}}$ in Eqn. (1) have disparate impacts on different demographic groups. Specifically, we consider two fairness metrics: group-wise loss disparity $\triangle_{\mathcal{L}}^{(t)}$ (martinez2020minimax; hashimoto2018fairness) and participation disparity $\triangle_{p}^{(t)}$ (Raab_Boczar_Fazel_Liu_2024), which measure the difference of group loss $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t)})$ and fraction $p_{s}^{(t)}$ across different groups at round $t$ of an iterative algorithm, respectively. In examples with two groups $\mathcal{S}=\{a,b\}$, the unfairness can be quantified as:
$\triangle_{\mathcal{L}}^{(t)}$ $:=\left|\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{a}^{(t)})-\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{b}^{(t)})\right|,$
$\triangle_{p}^{(t)}$ $:=\left|p_{a}^{(t)}-p_{b}^{(t)}\right|$ (4)
@@
Figure 1: Illustrating examples of polarization effects and unfairness of $\boldsymbol{\theta}^{\text{PS}}$ in Prop. 3.1 (left) and 3.2 (right): dynamics of the group fraction (left) and group-wise accuracy (right) under RRM: the system converges to $(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}})$ that is unfair (details are in App. D.2).
@@
We first show that PS solutions $\boldsymbol{\theta}^{\text{PS}}$ in SDPP may have disparate impacts on different groups. Specifically, when finding $\boldsymbol{\theta}^{\text{PS}}$ using iterative algorithms introduced in Section 2, the process may incur severe polarization effects and exhibit unfairness, i.e., certain groups may get more and more marginalized, and group-wise loss disparity gets exacerbated during the iterative process.
Proposition 3.1 (Polarization effects of $\boldsymbol{\theta}^{\text{PS}}$).
Consider population from multiple groups with fixed group distribution $\mathcal{D}_{s}$ whose participation $p_{s}^{(t+1)}$ in an ML system depends on their perceived group loss $\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s})$. Suppose the deployment of system $\boldsymbol{\theta}^{(t)}=\arg\min_{\boldsymbol{\theta}}\sum_{s\in\mathcal{S}}p_{s}^{(t)}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s})$ follows the conventional RRM schema, then there exist $\mathcal{D}^{(0)}$ and T such that as $t\to\infty$, $p_{s}^{(t)}$ changes monotonically and certain groups diminish entirely from the system.
Proposition 3.2 (Exacerbated group-wise loss disparity).
Consider population from multiple groups with fixed group proportion $p_{s}$; each individual is subject to a binary ML decision $\widehat{Y}^{(t)}=\mathbf{1}(X^{(t)}\geq\boldsymbol{\theta}^{(t)})$ and may strategically manipulate the data to increase the chance of receiving positive decisions. Suppose the deployment of ML system $\boldsymbol{\theta}^{(t)}=\arg\max_{\boldsymbol{\theta}}\Pr\big(\widehat{Y}^{(t)}=Y\big)$ follows the conventional RRM schema, and individuals manipulate features according to $X^{(t)}=X^{(t)}+\eta\boldsymbol{\theta}^{(t)}$ without changing YY(perdomo_performative_2021), then there exists $\mathcal{D}^{(0)}$ such that group-wise loss disparity $\Delta^{(t)}_{\mathcal{L}}$ increases.
To prove Prop. 3.1 and 3.2, it is sufficient to provide two examples to illustrate the polarization and unfairness effects in SDPP settings. We construct the examples with details in App. D.2 and visualize them in Figure 1. This shows that even though RRM can converge to a stable solution $\boldsymbol{\theta}^{\text{PS}}$, the solution is unfair and loss disparity $\triangle_{\mathcal{L}}^{(t)}$ and participation disparity $\triangle_{p}^{(t)}$ may get exacerbated.
4 Finding Fair-PS solutions
@@
Section 3 shows that without fairness consideration, PS solutions of SDPP may incur polarization effects and have disparate impacts on different groups. This section tackles unfairness issues in SDPP. One straightforward idea is to directly apply the fairness mechanisms at every round of the iterative algorithms introduced in Section 2. However, we will show that although such methods are effective in conventional supervised learning, they can disrupt the stability and the iterative algorithms may no longer converge.
4.1 Fair-PS solutions
@@
    Fairness via regularization: It adds a regularization or penalty term to the original learning objective function $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})$, which penalizes the violation of fairness (khan2023fairness; zhang2021unified). The fair objective function is
@@
    Fairness via sample re-weighting: It adjusts the weights of samples (possibly adversarially) and increases weights for disadvantaged groups (jung2023reweighting; duchi2018learning; duchi2023distributionally). An example is distributionally robust optimization (DRO) (pmlr-v80-hashimoto18a), which minimizes worst-case loss and the fair objective is
[…]
@@
$\mathcal{D}^{\text{PS}}_{\textsf{fair}}=\text{T}(\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}};\mathcal{D}^{\text{PS}}_{\textsf{fair}}),~~\boldsymbol{\theta}^{\text{PS}}_{\textsf{fair}}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}~\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{\text{PS}}_{\textsf{fair}},\rho).$
@@ 4.2 Existing designs fail to converge to Fair-PS solutions
We will use two examples to illustrate that the popular choices of fairness mechanisms used in existing literature may fail to achieve stability and fairness in SDPP. This includes (i) regularization method (khan2023fairness; zhang2021unified) with group loss variance $\mathcal{P}(\boldsymbol{\theta};\mathcal{D},\rho):=\rho\sum_{s\in\mathcal{S}}p_{s}\cdot[\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s})-\mathcal{L}(\boldsymbol{\theta};\mathcal{D})]^{2}$ as the penalty term in Eqn. (5); and (ii) distributionally robust optimization (DRO) method (hashimoto2018fairness; peetpare2023long) with $\chi^{2}$-distance metric $d(\mathcal{D},\tilde{\mathcal{D}}):=\int\left(\frac{d\mathcal{D}}{d\tilde{\mathcal{D}}}-1\right)^{2}d\tilde{\mathcal{D}}$ in Eqn. (6).
Example 4.2 (Group loss variance as penalty term).
Consider two groups $a,b$ with data $Z=(X,Y)$ and fixed group fractions $p_{a}=p_{b}=0.5$. Suppose all samples in group $a$ are (1,1)(1,-1) and all samples in group $b$ are $(2,1)$. Consider squared loss function $\ell(z;\theta)=(y-h_{\theta}(x))^{2}=(y-\theta x)^{2}$ which is strongly convex and jointly smooth. Under group loss variance penalty $\mathcal{P}(\boldsymbol{\theta};\mathcal{D},\rho)$, we have $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D},\rho)=2.5\theta^{2}-\theta+1+\rho\cdot(2.25\theta^{4}+9\theta^{2}-9\theta^{3})$. When $\rho=0.6$, the second-order derivative $\nabla^{2}_{\theta}\mathcal{L}_{\textsf{fair}}=16.2\theta^{2}-32.4\theta+15.8$ and is negative when $\theta=1$. The negative second-order gradient means that $\mathcal{L}_{\textsf{fair}}$ is nonconvex. According to Lemma 2.4, we cannot ensure the iterative algorithms converge to Fair-PS solutions when $\ell$ is nonconvex. Thus, adding group loss variance as a penalty at each round possibly disrupts the stability. Appendix C.5 verifies the non-convergence with empirical results.
Example 4.3 (Repeated DRO with $\chi^{2}$-distance metric).
@@
It is worth noting that DRO methods have been used in hashimoto2018fairness to mitigate group fraction disparity in repeated optimizations. However, it only improves fairness without any convergence guarantees to stable solutions. peetpare2023long repeatedly used DRO to improve fairness under PP settings, it only converges to a PS solution under stronger assumptions where the distributionally robust objective must be strongly convex and jointly smooth, and the transition map T also needs to be $\epsilon$-sensitive with respect to the worst-case distribution. Under milder conditions in Lemma 2.4, it may fail to converge as Example 4.3 illustrated.
@@ 4.3 Novel designs for fairness mechanism
Proposed fair regularization (with and without demographics).
Let $\mathcal{D}^{(0)}$ denote the initial population distribution. Depending on whether sensitive attributes $S$ are accessible during training, we propose two fairness penalty terms, as detailed below.
@@
    Group level fairness penalty: It updates $\boldsymbol{\theta}^{(t)}$ by minimizing $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ defined as follows
    $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})+\rho\sum_{s\in\mathcal{S}}p^{(t-1)}_{s}[\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})]^{2}$ (7)
@@
    Sample level fairness penalty without demographics: It updates $\boldsymbol{\theta}^{(t)}$ by minimizing $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ defined as follows, which does not require access to sensitive attribute values.
    $\textstyle\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})+\rho\mathbb{E}_{Z\sim\mathcal{D}^{(t-1)}}\big[[\ell(\boldsymbol{\theta};Z)]^{2}\big].$ (8)
Proposed fair sample re-weighting.
At the first round, it performs risk minimization. Starting from $t=2$, it updates $\boldsymbol{\theta}^{(t)}$ by minimizing $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ defined as follows
$\textstyle\sum_{s\in\mathcal{S}}q_{s}^{(t-1)}\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})$ (9)
with
$\textstyle\boldsymbol{q}^{(t-1)}$ $=\left[q^{(t-1)}_{s}\right]_{s\in\mathcal{S}}=\frac{\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}}{\|\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}\|_{1}}$
$\textstyle\boldsymbol{l}^{(t-1)}$ $=\left[p_{s}^{(t-1)}\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)})\right]_{s\in\mathcal{S}}$
Unlike DRO in peetpare2023long; hashimoto2018fairness that requires solving a min-max optimization at the current round, our re-weighting method only adjusts the weights for each group based on the group-wise losses in the previous round, which is more computationally efficient.
Comparison & discussion.
Intuitively, compared to original $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t-1)})$, all three proposed fair objective functions $\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$ improves fairness at each round by assigning more weights to disadvantaged groups/samples (i.e., those experiencing higher losses) in the upcoming update. Indeed, both sample level fairness penalty and fair sample re-weighting can be regarded as modifications of group level fairness penalty. Comparing Eqn. (7) and (8), the two penalty terms get similar when most individual samples in the disadvantaged (resp. advantaged) groups are also similarly disadvantaged (resp. advantaged). This is more likely to happen when each group has a distribution with a small variance. For example, if the distribution of each group is a point mass, Eqn. (7) and (8) are identical.
@@
$~~~~~~~\|\boldsymbol{p}^{(t-1)}+\rho\boldsymbol{l}^{(t-1)}\|_{1}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}^{(t-1)},\rho)$
$=\sum_{s}p_{s}^{(t-1)}(1+\rho\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)}))\,\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})$
We can see that the right-hand side will be a multiple of Eqn. (7) if we replace $\mathcal{L}(\boldsymbol{\theta}^{(t-1)};\mathcal{D}_{s}^{(t-2)})$ with $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s}^{(t-1)})$. This means both equations yield the same $\boldsymbol{\theta}^{(t)}$ when the population distribution does not change from t2t-2 to t1t-1, suggesting that the two approaches become more similar when the sensitivity of the transition map T is smaller, or equivalently, the distribution shift is milder.